The 4th term of an A.P. is equal to 3 times the first term and the 7th term exceeds wice the 3rd term by 1. Find the 1st term and d.

Grade:12th Pass

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

a+3d=3a

2a=3d

a+6d-a-2d=1

d=1/4

a=3/8